OpenStudy (anonymous):

am I right at all? Given that f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3. −15 << my answer −8 10 25

hero (hero):

@FrostFlare23 mind showing the work you did to arrive at that answer?

OpenStudy (anonymous):

alright give me a second

OpenStudy (anonymous):

well techincally I did work but none of the answers matched mine

OpenStudy (anonymous):

f(-3) = 2(-3) + 5

OpenStudy (anonymous):

f(-3) = -6 + 5 = -1

OpenStudy (anonymous):

did I screw up?

OpenStudy (anonymous):

@Hero ?

OpenStudy (anonymous):

you there dude?

hero (hero):

Looks right so far, but ...

hero (hero):

You should calculate f(g(x)) first, then calculate f(g(-3)).

OpenStudy (anonymous):

oh. uhm how do I do that? thats the only thing that confuses me

hero (hero):

I'll get you started: f(x) = 2x + 5 1. Replace x with g(x) f(g(x)) = 2(g(x)) + 5 2. Replace g(x) with x - 7 = 2(x - 7) + 5 3. Simplify the expression

OpenStudy (anonymous):

hmm... 2x -14 +5?

hero (hero):

So f(g(x)) = 2x - 14 + 5, but that can also be further reduced, right?

OpenStudy (anonymous):

I think so

hero (hero):

Because -14 + 5 = -9 so f(g(x)) = 2x - 9 Now calculate f(g(-3))

OpenStudy (anonymous):

hmm. so we would add 14 to both the 5 and the -14?

OpenStudy (anonymous):

oh sorry. lol late comment there

hero (hero):

@FrostFlare23 waiting on you ...

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so we would then divide 2 to both sides to isolate the X in order to get x = 4.5?

hero (hero):

No, you just simply evaluate f(g(x)) at x = -3 f(g(x)) = 2x - 9 f(g(-3)) = 2(-3) - 9

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so basically it would become: 2(-3) which equals -6. then we subtract -6 from 9 to get -15?

hero (hero):

Correct.

OpenStudy (anonymous):

ok so I was right I just ha the wrong way of doing it the first time thank you!

hero (hero):

Both ways of doing it are correct. It's just better practice to write f(g(x)) as a single expression then evaluate it once.

OpenStudy (anonymous):

alright thank you kind sir!

hero (hero):

The problem with doing it the way is, if someone asked you what f(g(x)) is, you wouldn't be able to tell them.

OpenStudy (anonymous):

oh ok